# Inflation via Moduli Potentials in a Nested Warped Geometry

### v1

@@ 4 Prospects for inflation
While it may be possible to solve the full system (4.4)--(4.7) numerically with appropriate initial conditions, we make here a simplifying assumption that would allow us to focus on the essential features of the multiply warped geometry at the cost of a reduced number of degrees of freedom. We assume one of the moduli to be chiefly responsible for driving inflation, with negligible rolling in the direction of the other modulus. In effect, this amounts to treating the latter modulus as a fixed parameter throughout an effective single-field inflationary epoch, which may be the result of a suitably implemented stabilization scheme. While this scenario may, at first glance, appear to be a little ad hoc, we remind the reader that the chief aim of the current work is not to propose a new model of non-canonical multi-field inflation per se, but rather to make a particular case study of one possible cosmological facet of curved multiply warped geometries. With this rationale in mind, we now proceed, with the aforementioned approximation, on to a case-by-case study of what each scenario implies for inflation and its observable features.
@@ 4.1 Inflation driven by $c$
Figure 4: Exact slow roll evolution of $c(N)$, $\epsilon(N)$, and $n_{s}(N)$ plotted for the initial condition $c(0)=2.419$, which results in $N_{\textrm{tot}}\approx 5420$ for $\omega_{2}=10^{-3}$. The evolution is independent of the value of the parameter $k$, i.e. trajectories for different plausible $k$-values (ranging between $10^{-3}$ and $10$) coincide. The smearing of $n_{s}(N)$ at smaller $e$-folds is a numerical artifact that gets resolved towards the tail end, which is of observational interest. This set of figures represents a viable scenario in the light of current CMB observations.
@@
The dynamics differ markedly depending on whether the field starts on the left or right of the inflection point. For $c(0)>c_{\textrm{in}}$, inflation does not come to an end through the violation of slow roll at any finite $N_{\textrm{tot}}$. This is because the non-canonical kinetic coefficient $\hat{\tilde{A}}_{cc}$, which results in a phantom-like kinetic term in this domain, drives the field up the potential slope towards the inflection point, with a gradually decreasing rolling speed and a subsequently decreasing value of $\epsilon(N)$. This behaviour of $\epsilon(N)$ can be understood from (4.9), where $\epsilon(N)$ is proportional to $\hat{V}_{,c}$ under the slow roll assumption. Since the latter approaches zero (by definition) as $c\to c_{\textrm{in}}$, $\epsilon(N)$ does not diverge and inflation can continue ad infinitum, unless some additional exit mechanism is invoked. This is displayed in Fig. 3 for a few benchmark values. On the other hand, for $c(0)<c_{\textrm{in}}$, inflation can come to an end naturally via slow roll violation. In this case, the phantom kinetic term makes $c(N)$ decrease over time in a similar manner as before, as the field is slowly driven up the potential plateau. However, $\hat{\tilde{A}}_{cc}$ changes its sign from positive to negative (for decreasing $c$) at a non-zero value of $c$ between $1$ and $c_{\textrm{in}}$, viz. upper right panel of Fig. 2. For $\omega_{2}=10^{-3}$, this critical value lies at $c\approx 2.1$, where $\hat{\tilde{A}}_{cc}\to 0$. As $\epsilon(N)$ is inversely proportional to $\hat{\tilde{A}}_{cc}$, this serves as the point where $|\epsilon(N)|$ diverges and inflation consequently ends, as shown in Fig. 4 for a case of physical interest. However, as this scenario involves momentary vanishing of the kinetic term, the analytic prescription adopted here may not be completely trustworthy close to this endpoint. To account for this, we conservatively allow an error margin of $\sim 0.2\%$ centered at the value of $N_{\textrm{tot}}$ where slow roll violation is naively expected. This poses little threat to our framework, as the total duration of inflation remains considerably larger than the minimum required amount.
@@
Table 1: For the CMB pivot scale $k_{*}=0.05$ Mpc-1, the interrelation among the inflationary quantities, i.e. the CMB exit $e$-fold $N_{k_{*}}$ from the beginning of inflation and the scalar spectral index $n_{s}(k_{*})$, and the braneworld parameters, i.e. the initial value $c(0)$, the parameter $k$, and the extra-dimensional radii ratio at horizon-crossing $(T_{1}/T_{2})_{*}$, has been shown for $c$-driven slow roll inflation. For values of $k$ outside this tabulated range, $(T_{1}/T_{2})_{*}$ quickly grows larger than $\mathcal{O}(10^{3})$, which amounts to a problematic hierarchy between the sizes of the extra dimensions. All the values correspond to $\omega_{2}=10^{-3}$ and $k^{\prime}\sim M$.
@@ 4.2 Inflation driven by $k$
Figure 5: Representative plots of $k(N)$, $\epsilon(N)$, and $n_{s}(N)$ for a few benchmark values of $c$, and a few particular initial condition imposed on $k$. For different initial conditions, the $k(N)$ curves run parallel to each other, leading eventually to different values of $N_{\rm tot}$ (upper left). For a given initial condition, the trajectories are independent of the value of $c$ and coincide with each other (upper right). The initial conditions chosen for the upper right panel result in exit from inflation around $N_{\rm tot}\sim 66-70$. But $n_{s}(N)$ turns out to be greater than unity throughout most of inflation, except very close to the exit where it drops sharply. All the plots correspond to $\omega_{2}=10^{-3}$ and $k^{\prime}\sim M$.
@@ 6 Discussions and conclusion
Following a brief recap of the essential geometric features of the curved doubly warped model in Sec. 2, we have started in Sec. 3 by explicitly deriving the 4D induced potential and the modular kinetic terms in the Jordan frame, which involves a non-minimal coupling of the 4D Ricci scalar to the moduli fields themselves. This has been achieved by splitting the total 6D Ricci scalar appropriately and identifying the specific type of effective lower-dimensional term arising from each part. The derivatives with respect to the extra-dimensional angular coordinates, alongside the bulk cosmological constant and the 4-brane tension terms in the full 6D action, lead to the effective potential $V(c,k)$, which shows distinct behaviour in the dS and AdS regimes (Sec. 3.1). The derivatives with respect to the usual non-compact 4D coordinates, on the other hand, lead to the effective kinetic terms after suitable integration by parts, whereas the non-minimal Jordan frame coupling emerges from the fact that the 6D Ricci scalar contains the warped 4D Ricci scalar within itself (Sec. 3.2). The entire Jordan frame action has then been recast in the Einstein frame in Sec. 3.3, where the non-minimal gravitational coupling is removed at the cost of non-trivial modifications (involving the same coupling function) to the form of the effective potential, as well as to those of the kinetic terms. For $\Omega>0$ in particular, which is the case of physical interest, the Einstein frame potential $\hat{V}(c,k)$ contains a metastable plateau of finite width in the direction of the first modulus $c$, while decaying with no minima along the second modulus $k$. Thus, it appears that positively curved brane geometry alone cannot stabilize all of the moduli in a metastable fashion in multiply warped settings, and it becomes important to invoke additional stabilizing mechanisms for the higher moduli. On the other hand, the non-canonical kinetic couplings in the Einstein frame are found to depend on the moduli following a nice pattern, with the coupling function of each self-kinetic term depending only on the corresponding modulus, and that of the single cross-kinetic term depending on both $c$ and $k$. This is but a consequence of the nested geometric structure, which allows such a splitting in the Einstein frame which is equipped with a minimal gravitational coupling. Moreover, the input of the brane curvature appears in only those kinetic terms which involve at least one derivative of $c(x)$. This, in turn, is a consequence of the fact that only the first warp factor $a(y)$ (which is associated with $c$) is explicitly dependent on $\Omega$, which does not show up in the expression of any subsequently higher warp factor. As we demonstrate by means of a 7D extension in Sec. 5, most of these dynamical features can be generalized readily to any arbitrary $n$-fold warped braneworld with $n\geq 2$ in presence of non-zero brane curvature.
@@
In Sec. 4, we have made use of the results derived in Sec. 3 to study the prospects for slow roll inflation, which exemplifies a physical scenario where the idea of such a curved braneworld may prove particularly useful. While a full-fledged two-field analysis involving both of the moduli is possible in principle, it is numerically expensive and not entirely relevant within the framework of the present study, which primarily aims to investigate, as part of its full scope, some possible cosmological phenomenon which might be addressed naturally from the perspective of a non-flat multiply warped braneworld. As such, it would be interesting to pursue the former route in future as part of a detailed study focused purely on braneworld-induced multi-field inflation. In this work, we have restricted ourselves to a case-by-case analysis instead, where we have explored the consequences of slow roll along the potential predominantly in one particular modular direction at a time, with the other modulus held approximately fixed and thus playing a parametric role. We maintain an agnostic stance towards the process(es) which may stabilize the parametric modulus, and focus only on the possible observational consequences of such a scenario. Nonetheless, in fundamentally string-inspired frameworks which may give rise to effective warped geometries in lower dimensions, it is not altogether impossible to imagine such moduli stabilizing processes resulting from $p$-form fluxes on the higher-dimensional Calabi-Yau manifolds [82, 83, 84, 85]. We do not pursue this line of inquiry in the present work any further, and leave it to a future study.
@@
Proceeding in the spirit described above, we have investigated inflationary scenarios driven individually by $c(x)$ and $k(x)$ in Sec. 4.1 and Sec. 4.2 respectively. Firstly, it is interesting to note that both of the cases come equipped with in-built exit mechanisms at some finite $e$-fold value, which is linked to the nature of the non-canonical kinetic coefficients $\hat{\tilde{A}}_{cc}(c)$ and $\hat{\tilde{A}}_{kk}(k)$. Each of these coefficients exhibits a zero-crossing in its corresponding field space, where the kinetic term switches its sign. For suitable initial values greater than this critical value, the kinetic term is phantom, and the field slowly climbs up the potential towards this tipping point. The slow roll parameter, being inversely proportional to the non-canonical kinetic coefficient, eventually diverges at this point, marking an exit from inflation. We would like to emphasize that our scenario is distinct from other inflationary proposals involving negative kinetic terms, e.g. phantom inflation [86, 87, 88] or ghost inflation [89, 90, 91], due to the simultaneous presence of the non-trivial kinetic coefficients and the effective potential, neither of which needs to be invoked ad hoc in our case.
@@
In the end, we would like to note that the present analysis is by no means exhaustive, and that there may be numerous interesting extensions. Reiterating our comment from earlier, the first avenue to explore could be whether it is possible to obtain a concrete stabilization mechanism which may justify the unidirectional slow roll scheme and the stability of the other moduli, which we have explored as particular cases of interest in this analysis. This could be studied, for instance, in the context of Type IIB string-inspired frameworks where fluxes play an important stabilizing role. The second avenue would be to proceed with a full numerical computation of both the background and perturbations by keeping all the moduli free, thus approaching the scenario from the perspective of multi-field inflation which we have deliberately avoided in this work. Apart from these two, other possibilities include studying late time cosmological effects in the dS branch, or potential physical applications of the AdS branch, which have been relatively less explored even in the 5D scenario. We plan to report on some of these aspects in future works.
@@ Appendix B Plots of slope-to-height ratio of the potential
The nature of $(\hat{V}_{,c}/\hat{V})$ and $(\hat{V}_{,k}/\hat{V})$, which appear respectively in (4.8) and (4.12), are shown graphically as follows.

### v2

@@ 4 Prospects for inflation
While it may be possible to solve the full system (4.4)--(4.7) numerically with appropriate initial conditions, we make here a simplifying assumption that would allow us to focus on the essential features of the multiply warped geometry at the cost of a reduced number of degrees of freedom. We assume one of the moduli to be chiefly responsible for driving inflation, with negligible rolling in the direction of the other modulus. In effect, this amounts to treating the latter modulus as a fixed parameter throughout an effective single-field inflationary epoch, which may be the result of a suitably implemented stabilization scheme. While this scenario may, at first glance, appear to be a little ad hoc, we remind the reader that the chief aim of the current work is not to propose a new model of non-canonical multi-field inflation per se, but rather to make a particular case study of one possible cosmological facet of curved multiply warped geometries. That being said, even within the present framework alone, it is possible to physically motivate the assumed unidirectional nature of the slow-roll based on certain quantitative arguments which we discuss in Appendix C. With this rationale in mind, we now proceed, with the aforementioned approximation, on to a case-by-case study of what each scenario implies for inflation and its observable features.
@@ 4.1 Inflation driven by $c$
Figure 5: Exact slow roll evolution of $c(N)$, $\epsilon(N)$, and $n_{s}(N)$ have been plotted for the initial condition $c(0)=2.419$, which results in $N_{\textrm{tot}}\approx 5420$ for $\omega_{2}=10^{-3}$. The evolution is independent of the value of the parameter $k$, i.e. trajectories for different plausible $k$-values (ranging between $10^{-3}$ and $10$) coincide. Note that in the upper-right plot, inflation proceeds from right to left, and the steep upward curve towards the left margin actually represents the divergence of $|\epsilon(N)|$ close to that point signifying the the end of inflation. While an extended vertical plot range better illustrates this divergence, it visually flattens the rest of the curve and makes it difficult to show the precise evolution of $\epsilon(N)$ throughout the course of inflation. The smearing of $n_{s}(N)$ in the bottom plot at smaller $e$-folds is a numerical artifact that gets resolved towards the tail end, which is of observational interest. This set of figures represents a viable scenario in the light of current CMB observations.
@@
The dynamics differ markedly depending on whether the field starts on the left or right of the inflection point. For $c(0)>c_{\textrm{in}}$, inflation does not come to an end through the violation of slow roll at any finite $N_{\textrm{tot}}$. This is because the non-canonical kinetic coefficient $\hat{\tilde{A}}_{cc}$, which results in a phantom-like kinetic term in this domain, drives the field up the potential slope towards the inflection point, with a gradually decreasing rolling speed and a subsequently decreasing value of $\epsilon(N)$. This behaviour of $\epsilon(N)$ can be understood from (4.9), where $\epsilon(N)$ is proportional to $\hat{V}_{,c}$ under the slow roll assumption. Since the latter approaches zero (by definition) as $c\to c_{\textrm{in}}$, $\epsilon(N)$ does not diverge and inflation can continue ad infinitum, unless some additional exit mechanism is invoked. This is displayed in Fig. 4 for a few benchmark values. On the other hand, for $c(0)<c_{\textrm{in}}$, inflation can come to an end naturally via slow roll violation. In this case, the phantom kinetic term makes $c(N)$ decrease over time in a similar manner as before, as the field is slowly driven up the potential plateau. However, $\hat{\tilde{A}}_{cc}$ changes its sign from positive to negative (for decreasing $c$) at a non-zero value of $c$ between $1$ and $c_{\textrm{in}}$, viz. upper right panel of Fig. 2. For $\omega_{2}=10^{-3}$, this critical value lies at $c\approx 2.1$, where $\hat{\tilde{A}}_{cc}\to 0$. As $\epsilon(N)$ is inversely proportional to $\hat{\tilde{A}}_{cc}$, this serves as the point where $|\epsilon(N)|$ diverges and inflation consequently ends, as shown in Fig. 5 for a case of physical interest. The upper-right plot of Fig. 5, however, does not visually reflect this divergence to a high level of accuracy, and shows only a steep upward curve close to the end (left margin) instead. This upward curve actually represents the aforesaid divergence, which is hard to depict graphically in a clearer manner. While extending the vertical range of the plot can better show the divergence, extending it to $\mathcal{O}(1)$ (as required to concretely show the end of inflation) visually flattens the rest of the curve and makes the plot lose information about precisely how $\epsilon(N)$ varies throughout inflation. Hence, we present the $N$ versus $\epsilon(N)$ plot in the form that appears in Fig. 5, keeping in mind that a divergence of the slow-roll parameter actually ends the inflating scenario as indicated by the steep upward rise of $|\epsilon(N)|$ close to $N_{\rm tot}\sim 5420$. Just to clarify at this point once more, recall that the range of $e$-folds of observable interest in cosmology consists almost entirely of the final $50-60$ $e$-foldings before the end of inflation, which, in case of Fig. 5, is comprised of $N\sim 5360-5370$. However, we have plotted the evolution of $c(N)$, $\epsilon(N)$, and $n_{s}(N)$ throughout the entire course of inflation (which is markedly larger in our case than the physically observable range), in order to clearly illustrate the theoretical nature of the trajectories as well as to make it easier to connect Fig. 5 with the initial values chosen in our calculations. Moreover, as this scenario involves momentary vanishing of the kinetic term, the analytic prescription adopted here may not be completely trustworthy close to this endpoint. To account for this, we conservatively allow an error margin of $\sim 0.2\%$ centered at the value of $N_{\textrm{tot}}$ where slow roll violation is naively expected. This poses little threat to our framework, as the total duration of inflation remains considerably larger than the minimum required amount.
@@
Table 1: For the CMB pivot scale $k_{*}=0.05$ Mpc-1, the interrelation among the inflationary quantities, i.e. the CMB exit $e$-fold $N_{k_{*}}$ from the beginning of inflation and the scalar spectral index $n_{s}(k_{*})$, and the braneworld parameters, i.e. the initial value $c(0)$, the parameter $k$, and the extra-dimensional radii ratio at horizon-crossing $(T_{1}/T_{2})_{*}$, has been shown for $c$-driven slow roll inflation. For values of $k$ outside this tabulated range, $(T_{1}/T_{2})_{*}$ quickly grows larger than $\mathcal{O}(10^{3})$, which amounts to a problematic hierarchy between the sizes of the extra dimensions. Recall that $\hat{V}(c,k)$ is dimensionless by definition, and all the values correspond to $\omega_{2}=10^{-3}$ and $k^{\prime}\sim M$.
@@ 4.2 Inflation driven by $k$
Figure 6: Representative plots of $k(N)$, $\epsilon(N)$, and $n_{s}(N)$ for a few benchmark values of $c$, and a few particular initial condition imposed on $k$. For different initial conditions, the $k(N)$ curves run parallel to each other, leading eventually to different values of $N_{\rm tot}$ (upper left). For a given initial condition, the trajectories are independent of the value of $c$ and coincide with each other (upper right). The initial conditions chosen for the upper right panel result in exit from inflation around $N_{\rm tot}\sim 66-70$. But $n_{s}(N)$ turns out to be greater than unity throughout most of inflation, except very close to the exit where it drops sharply. Note that in the bottom-left plot, inflation proceeds from right to left, and the steep upward curve towards the left margin actually represents the divergence of $|\epsilon(N)|$ close to that point signifying the the end of inflation. While extending the vertical range of the plot might better illustrate this divergence, it visually flattens the rest of the curve and makes it difficult to show the precise evolution of $\epsilon(N)$ throughout the course of inflation. All the plots correspond to $\omega_{2}=10^{-3}$ and $k^{\prime}\sim M$.
@@ 6 Discussions and conclusion
Following a brief recap of the essential geometric features of the curved doubly warped model in Sec. 2, we have started in Sec. 3 by explicitly deriving the 4D induced potential and the modular kinetic terms in the Jordan frame, which involves a non-minimal coupling of the 4D Ricci scalar to the moduli fields themselves. This has been achieved by splitting the total 6D Ricci scalar appropriately and identifying the specific type of effective lower-dimensional term arising from each part. The derivatives with respect to the extra-dimensional angular coordinates, alongside the bulk cosmological constant and the 4-brane tension terms in the full 6D action, lead to the effective potential $V(c,k)$, which shows distinct behaviour in the dS and AdS regimes (Sec. 3.1). The derivatives with respect to the usual non-compact 4D coordinates, on the other hand, lead to the effective kinetic terms after suitable integration by parts, whereas the non-minimal Jordan frame coupling emerges from the fact that the 6D Ricci scalar contains the warped 4D Ricci scalar within itself (Sec. 3.2). The entire Jordan frame action has then been recast in the Einstein frame in Sec. 3.3, where the non-minimal gravitational coupling is removed at the cost of non-trivial modifications (involving the same coupling function) to the form of the effective potential, as well as to those of the kinetic terms. For $\Omega>0$ in particular, which is the case of physical interest, the Einstein frame potential $\hat{V}(c,k)$ contains a metastable plateau of finite width in the direction of the first modulus $c$, while decaying with no minima along the second modulus $k$. Thus, it appears that positively curved brane geometry alone cannot stabilize all of the moduli in a metastable fashion in multiply warped settings, and it becomes important to invoke additional stabilizing mechanisms for the higher moduli. On the other hand, the non-canonical kinetic couplings in the Einstein frame are found to depend on the moduli following a nice pattern, with the coupling function of each self-kinetic term depending only on the corresponding modulus, and that of the single cross-kinetic term depending on both $c$ and $k$. This is but a consequence of the nested geometric structure, which allows such a splitting in the Einstein frame which is equipped with a minimal gravitational coupling. Moreover, the input of the brane curvature appears in only those kinetic terms which involve at least one derivative of $c(x)$. This, in turn, is a consequence of the fact that only the first warp factor $a(y)$ (which is associated with $c$) is explicitly dependent on $\Omega$, which does not show up in the expression of any subsequently higher warp factor. As we demonstrate by means of a 7D extension in Sec. 5, most of these dynamical features can be generalized readily to any arbitrary $n$-fold warped braneworld with $n\geq 2$ in presence of non-zero brane curvature. Furthermore, in Sec. 3.4, we have explicitly computed the field space curvature for the non-canonical kinetic scenario that emerges in the Einstein frame, in both the AdS and the dS regimes separately. This analysis supplements the preceding calculations in a more robust manner, and illustrates interesting trends in the variation of non-canonicity as a function of the dimensionless moduli fields $c$ and $k$.
@@
In Sec. 4, we have made use of the results derived in Sec. 3 to study the prospects for slow roll inflation, which exemplifies a physical scenario where the idea of such a curved braneworld may prove particularly useful. While a full-fledged two-field analysis involving both of the moduli is possible in principle, it is numerically expensive and not entirely relevant within the framework of the present study, which primarily aims to investigate, as part of its full scope, some possible cosmological phenomenon which might be addressed naturally from the perspective of a non-flat multiply warped braneworld. As such, it would be interesting to pursue the former route in future as part of a detailed study focused purely on braneworld-induced multi-field inflation. In this work, we have restricted ourselves to a case-by-case analysis instead, where we have explored the consequences of slow roll along the potential predominantly in one particular modular direction at a time, with the other modulus held approximately fixed and thus playing a parametric role. While we maintain an agnostic stance towards any possible process(es) which may absolutely stabilize the parametric modulus, we find that it is nonetheless possible to achieve a nearly constant value of any one modulus field while allowing the other one to roll, in specific regions of the field space in our scenario. Besides, in fundamentally string-inspired frameworks which may give rise to effective warped geometries in lower dimensions, it is not altogether impossible to imagine more sophisticated moduli stabilizing processes resulting from $p$-form fluxes on the higher-dimensional Calabi-Yau manifolds [84, 85, 86, 87]. We have not pursued this line of inquiry any further and left it to a future study, focusing instead on the physically observable consequences of such unidirectional slow-roll scenarios.
@@
To that end, we have investigated inflationary scenarios driven individually by $c(x)$ and $k(x)$ in Sec. 4.1 and Sec. 4.2 respectively. Firstly, it is interesting to note that both of the cases come equipped with in-built exit mechanisms at some finite $e$-fold value, which is linked to the nature of the non-canonical kinetic coefficients $\hat{\tilde{A}}_{cc}(c)$ and $\hat{\tilde{A}}_{kk}(k)$. Each of these coefficients exhibits a zero-crossing in its corresponding field space, where the kinetic term switches its sign. For suitable initial values greater than this critical value, the kinetic term is phantom, and the field slowly climbs up the potential towards this tipping point. The slow roll parameter, being inversely proportional to the non-canonical kinetic coefficient, eventually diverges at this point, marking an exit from inflation. We would like to emphasize that our scenario is distinct from other inflationary proposals involving negative kinetic terms, e.g. phantom inflation [88, 89, 90] or ghost inflation [91, 92, 93], due to the simultaneous presence of the non-trivial kinetic coefficients and the effective potential, neither of which needs to be invoked ad hoc in our case.
@@
In the end, we would like to note that the present analysis is by no means exhaustive, and that there may be numerous interesting extensions. Reiterating our comment from earlier, the first avenue to explore could be whether it is possible to obtain a concrete stabilization mechanism which may justify the unidirectional slow roll scheme and the stability of the other moduli more rigorously, which we have explored as particular cases of interest in this analysis. This could be studied, for instance, in the context of Type IIB string-inspired frameworks where fluxes play an important stabilizing role. The second avenue would be to proceed with a full numerical computation of both the background and perturbations by keeping all the moduli free, thus approaching the scenario from the perspective of multi-field inflation which we have deliberately avoided in this work. Apart from these two, other possibilities include studying late time cosmological effects in the dS branch, or potential physical applications of the AdS branch, which have been relatively less explored even in the 5D scenario. We plan to report on some of these aspects in future works.
@@ Appendix B Plots of slope-to-height ratio of the potential
The nature of $(\hat{V}_{,c}/\hat{V})$ and $(\hat{V}_{,k}/\hat{V})$, which appear respectively in (4.8) and (4.12), has been shown graphically in Fig. 8.
@@
Appendix C Physical intuition behind unidirectional slow-roll
In this appendix, we briefly discuss the physical reasoning behind our consideration of slow-roll along either field direction at a time, with the other field value held approximately constant. First, we focus on slow-roll along the $c$-direction as considered in Sec. 4.1, for which the fixed nature of $k$ might be motivated in a step-by-step manner as follows.
  1. 1.
    In Fig. 2, we have plotted the potential and the non-canonical coefficients in the dS regime. In Fig. 2, focusing on the region $c\in(2,4)$, we see that the cross-coefficient $\hat{\tilde{A}}_{ck}(c,k)$ becomes almost flat along both $c$ and $k$ for all values of $k$, thus rendering both $\hat{\tilde{A}}_{ck,c}$ and $\hat{\tilde{A}}_{ck,k}$ negligibly small.
  2. 2.
    On the other hand, the value of $\hat{\tilde{A}}_{ck}$ is of the typical order $-\mathcal{O}(10)$ in this region.
  3. 3.
    Next, as (4.6) and (4.7) are second order ODEs, we can freely choose a vanishing initial condition $\dot{k}_{\rm in}\to 0$ for the rolling speed.
  4. 4.
    Implementing the three pieces of information above in (4.6) and (4.7), the initial stage of the coupled dynamics may be well approximated by
    $2\hat{\tilde{A}}_{cc}(\ddot{c}+3H\dot{c})+\hat{\tilde{A}}_{ck}\ddot{k}-2k^{\prime 2}\hat{V}_{,c}\approx 0\>,$ (C.1)
    $2\hat{\tilde{A}}_{kk}\ddot{k}+\hat{\tilde{A}}_{ck}(\ddot{c}+3H\dot{c})-2k^{\prime 2}\hat{V}_{,k}\approx 0\>.$ (C.2)
  5. 5.
    Now, for $k\sim 1.0$ and $c$ starting close to the inflection point $c_{\rm in}\approx 2.41994$, the typical magnitudes of the three non-canonical coefficients in the equations above turn out to be $\hat{\tilde{A}}_{cc}\sim 10^{-4}$, $\hat{\tilde{A}}_{kk}\sim 321$, and $\hat{\tilde{A}}_{ck}\sim-15$. Plugging these values in (C.2), we get the order-of-magnitude estimate
    $\ddot{c}+3H\dot{c}\sim(15\ddot{k}+2k^{\prime 2}\hat{V}_{,c})\times 10^{4}\>.$ (C.3)
    Using this in (C.1), we get an order-of-magnitude estimate for $\ddot{k}$ as
    $\ddot{k}\sim-k^{\prime 2}(30\hat{V}_{,c}+2\hat{V}_{,k})\times 10^{-6}\>.$ (C.4)
  6. 6.
    Referring to Fig. 2, we see that the slopes of the potential $\hat{V}(c,k)$ along both $c$ and $k$ become very small in the region mentioned above in (e), i.e., both $|\hat{V}_{,c}|$ and $|\hat{V}_{,k}|$ are $\ll\mathcal{O}(1)$. In fact, a direct computation reveals $|\hat{V}_{,c}|\sim\mathcal{O}(10^{-15})$ and $|\hat{V}_{,k}|\sim\mathcal{O}(10^{-8})$. So, overall, $|\ddot{k}|\ll\mathcal{O}(k^{\prime 2})$. This, together with the previously chosen initial condition $\dot{k}_{in}\to 0$, essentially makes the time-variation of the $k$-field negligibly small, which effectively leads us to the $c$-driven case which is the focus of Sec. 4.1.
Finally, while the arguments above have been presented for $k\sim 1.0$ for the purpose of demonstration, it is easy to see that any change in the starting value of $k$ affects only the two quantities $\hat{\tilde{A}}_{kk}$ and $\hat{V}_{,k}$. For $k\to 0$, both $\hat{A}_{kk}\to-\infty$ and $\hat{V}_{,k}\to-\infty$, as observed in Figs. 2 and 2. But for any positive value of $k$ that is sufficiently away from this divergence (e.g. $k\gtrsim 0.1$), the order of magnitude calculations presented above remain valid and render $\ddot{k}$ quite small. This, in turn, justifies most of the $k$-values considered in Table 1, with the possible exception of the two smallest values. However, we have included those values in our analysis from an inclusive standpoint, in case they too might be motivated physically in the future in extended frameworks.
In a similar spirit, the fixed nature of $c$ in case of $k$-driven slow-roll may be demonstrated. Within the range $c\sim 0.1-1.0$ and $k\sim 7.0-11.0$ as chosen in Sec. 4.2, choosing $\dot{c}_{\rm in}\to 0$ allows us to approximate the initial dynamics described by (4.6) and (4.7) by
$2\hat{\tilde{A}}_{kk}(\ddot{k}+3H\dot{k})+\hat{\tilde{A}}_{ck}\ddot{c}-2k^{\prime 2}V_{,k}\approx 0\>,$ (C.5)
$2\hat{\tilde{A}}_{cc}\ddot{c}+\hat{\tilde{A}}_{ck}(\ddot{k}+3H\dot{k})-2k^{\prime 2}V_{,c}\approx 0\>.$ (C.6)
If, say, $c\sim 0.5$ and $k\sim 7.0$ (for demonstration), then the typical magnitudes of the quantities involved are $\hat{\tilde{A}}_{kk}\sim 325$, $\hat{\tilde{A}}_{cc}\sim-6$, and $\hat{\tilde{A}}_{ck}\sim-60$. Using the first equation to eliminate $(\ddot{k}+3H\dot{k})$ in the second equation, the acceleration of the $c$-field is of the order
$\ddot{c}\sim 0.1\times k^{\prime 2}(\hat{V}_{,c}+0.1\times\hat{V}_{,k})\>.$ (C.7)
In this field zone of interest, we obtain $|\hat{V}_{,c}|\sim\mathcal{O}(10^{-34})$ and $|\hat{V}_{,k}|\sim\mathcal{O}(10^{-32})$, which are vanishingly small. Thus, $\ddot{c}$ is negligibly small, which, together with the initial speed $\dot{c}_{\rm in}\to 0$, effectively freezes the $c$-field and reduces the scenario to the one considered in Sec. 4.2.